OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

jigglypuff314 (jigglypuff314):

216 is 648/3 72 is 216/3 24 is 72/3 etc

OpenStudy (mathstudent55):

Test the sequence to see if it's an arithmetic sequence. Subtract each term from the next term. 216 - 648 = 432 72 - 216 = 144 Since the differences are not the same, there is no common difference. That means the sequence is not an arithmetic sequence. Now test the sequence for a geometric sequence. Divide each term by the previous term. 24/72 = 1/3 72/216 = 1/3 216/648 = 1/3 Since all the divisions give the same result, there is a common factor between terms. That means it's a geometric sequence. Each term is one-third of the previous term. Now you can use that to find the next three terms.